# Collaborative filtering based on nonnegative/binary matrix factorization

### v2

@@ Methods
The method to extract data at a specified filling rate is as follows. First, we sorted the columns in the descending order of the percentage of filled elements in each column and then sorted the rows in a similar manner. Next, we selected an $n\times m$ matrix whose $(1,1)$ element coincides with the first-row and first-column element of the sorted table, and calculated the filling rate of the matrix. By shifting the $(1,1)$-element location by one row and one column in the sorted table, we repeated the calculation of the filling rate. The $n\times m$ matrix whose filling rate was closest to the desired filling rate was selected as the rating matrix with the desired filling rate.
Ising formulation
$E=-\frac{1}{2}\sum_{i,j}J_{ij}s_{i}s_{j}+\sum_{i}h_{i}s_{j}.$ (15)
To minimize Eq. (9) to update the $j$th row of $M\circ(WH)$, $W_{rl}$ in Eqs. (16) and (17) are replaced with $(\tilde{W})_{rl}=M_{rj}W_{rl}$.
Simulated-bifurcation-based Ising machine
The SB algorithm, which is based on the adiabatic evolution in classical nonlinear systems that exhibit bifurcation, was introduced to accelerate combinatorial optimization [23, 24, 25]. The SB algorithm has several variants, including adiabatic, ballistic, and discrete SB. In this study, we employed the ballistic SB method, whose update rule is described below [25].
$x_{i}(t_{k+1})$ $=x_{i}(t_{k})+a_{0}y_{i}(t_{k+1})\Delta_{t},$ (19)
In our experiments, we employed a device with the FPGA implementation of the SB algorithm (SB-based Ising machine), the same one as in Ref. 26, to minimize Eqs. (5) and  (9). The SB-based Ising machine (Figure 7) can solve fully-connected 2,048-spin Ising problems (the machine size $M$ is 2,048), featuring a computational precision of 32-bit floating points and a system clock frequency of 259 MHz. As shown in Figure 7a, the FPGA (Intel Stratix 10 SX 2800 FPGA) on the board (Intel FPGA PAC D5005 accelerator card) is connected to a CPU (Intel Core i9-9900K, 3.60 GHz) via a PCIe bus (Gen 3$\times$16, peak bandwidth of 15.75 GB/s). The NBMF process is executed using a CPU; however, the Ising problems described in Eqs. (16) and (17) are transferred/solved (offloaded) to/using the SB-based Ising machine. The computation times shown in Figures 1, 3, and 5 include the processing times of the CPU and FPGA and the data transfer times (overhead times) between them. The NMF process was executed only on the CPU (no data transfer time). The column update problems involved in updating the matrix $H$ [Eq. (5)], each formulated as an Ising problem of size $k$ [Eqs. (16) and (17)], are independent and thus can be processed simultaneously. By packing the multiple-column update problems as a large Ising problem, as shown in Figure 7b [placing the small $J$ matrices on the diagonal line with zero padding to the remaining off-diagonal components], we solve $\lfloor M/k\rfloor$ column update problems simultaneously using the SB-based Ising machine with size $M$, where $\lfloor A\rfloor$ is the floor function of $A\in\mathbb{R}$.
Data Availability
Data supporting the findings of this study are available from the corresponding author upon request.
Acknowledgements
This study was partially supported by JSPS KAKENHI (Grant Number JP23H04499) and Murata Science and Education Foundation.
Author contributions statement
@@ Acknowledgements
Additional information
Competing interests: K.T. is an inventor on a Japanese patent application related to this work filed by the Toshiba Corporation (no. P2019-164742, filed 10 September 2019). This study was conducted as collaborative research between Ochanomizu University and Toshiba Corporation. The authors declare that they have no other competing interests.

### v3

@@ III.6 Data preparation
The method for extracting data at a specified filling rate is as follows. First, we sorted the columns in descending order by the percentage of filled elements in each column and then sorted the rows similarly. Next, we selected an $n\times m$ matrix whose $(1,1)$ element coincides with the first-row and first-column element of the sorted table, and calculated the filling rate of the matrix. By shifting the $(1,1)$-element location by one row and one column in the sorted table, we repeated the calculation of the filling rate. The $n\times m$ matrix with the closest filling rate to the desired filling rate was selected as the rating matrix.
III.7 Parameter settings and evaluation
By extracting data from each dataset, we constructed a rating matrix in which 20% of the elements were rated unless otherwise specified. The numbers of users (rows) and items (columns) in the rating matrix are $n=250$ and $m=500$, respectively, and the number of features is set to half the number of users, that is, $k=n/2$, unless otherwise specified. For the learning process, which involved the execution of NBMF/NMF, we concealed 20% of the rated elements together with the unrated ones. To evaluate the performance, we used the root mean squared error (RMSE) of the rated elements:
We set the hyperparameters in Eqs. (2) and (3) as $\lambda_{1}=10^{-2}$ and $\lambda_{2}=10^{-5}$, which were tuned for the MovieLens dataset using a grid search. Parameters in Eqs. (9) and (17) were also tuned for the case of the MovieLens dataset. Although the optimal values may depend on the dataset and matrix size, we used the fixed values for simplicity.
In our experiments, we applied NBMF and NMF to the same rating matrix. To ensure a comprehensive evaluation, we divided the rated elements into five distinct sets and performed five trials, masking one set at a time. The average was calculated for five trials unless otherwise specified.
IV Results and discussion
Figure 3 shows a comparison of RMSE and computation time of NBMF and NMF for 10 epochs. Each epoch involves updating matrix $W$ followed by updating matrix $H$. The data points represent the averages of RMSE and computation time at each epoch, with some error bars too small to be observed. Figure 3 shows that the RMSE decays more rapidly in NBMF than in NMF for all datasets. Although the difference in the RMSE at each epoch between NBMF and NMF was negligible, the difference among the datasets was remarkable. The difference among the datasets originates from the frequency distribution of the ratings in each dataset, as shown in Figure 6. As elaborated later, when the distribution was sharp and the variance was small, the RMSE tended to be small. In contrast, when the variance in the frequency was large, the RMSE was relatively large.
Figure 4: Filling rate dependence of RMSE for (a) MovieLens, (b) Netflix, (c) Yahoo, and (d) CiaoDVD datasets. Training data with filling rates of more than 45%, 25%, and 20% could not be extracted for (b), (c), and (d), respectively. The data were averaged over five trials, with error bars denoting the standard deviation.
The filling rate of a rating matrix, which is the proportion of rated elements, influences collaborative filtering. However, the filling rate dependence of the RMSE varied across the datasets, as shown in Figure 4. Here, the RMSE was calculated after 10 epochs and averaged over five trials.
NMF is is expected to produce lower RMSE values than NBMF due to its higher resolution. However, in Figure 4, case (a) demonstrates that the RMSE values for both NBMF and NMF were similar when the filling rate was around 20%. Furthermore, in case (d), the RMSE for NBMF was smaller than that for NMF at the filling rate of approximately 20%. This inconsistent behavior suggests that the datasets have significant differences in their features.
Figure 5 shows the advantages of this masking procedure. The masking data for each dataset were the same as those used for NBMF in Figure 3. Notably, the RMSE without the masking procedure was about more than three times larger than the RMSE with the masking procedure for all datasets. Furthermore, the difference in computation time between the two procedures was minimal. These findings indicate that the masking procedure provides apparent benefits in collaborative filtering, despite a minor drawback in terms of computation time.
The results indicate that the RMSE reflects specific properties of the data. Here, we focus on the frequency distribution of ratings, as illustrated in Figure 6. The distribution represents the percentage of ratings (1, 2, 3, 4, and 5) among the rated elements in a rating matrix, with a filling rate of 20%. The distributions in (a) and (b) showed a broad peak, and the corresponding RMSE had a similar value at 10 epochs in Figures 3a and 3b. However, the distribution in (c) showed two peaks at 1 and 5, resulting in significant variance. The corresponding RMSE at 10 epochs in Figure 3c was larger than those of the other three data. By contrast, the distribution in (d) had a steep peak at 4, indicating that more than 80% of the rated elements had a value of 4. The corresponding RMSE at 10 epochs in Figure 3d was significantly smaller than those of the other three data. This observation indicates that a distribution with a sharp peak and small variance typically results in a smaller RMSE. However, a distribution with a broader peak and larger variance often results in a larger RMSE.
The computation time for NBMF was significantly shorter than that for NMF under the same problem setup, as shown in Figure 7, for all the datasets. The total time required to update matrix $W$ over 10 epochs was the same for both NMF and NBMF. However, the time required to update matrix $H$ in NMF was approximately six times longer than that in NBMF. This discrepancy suggests that the use of an SB-based machine accelerates the computation to update $H$. Additional time is required to minimize Eq. (13) during the update of $H$. Executing minimization using the SB-based machine involves transforming the objective function into the Ising model form, as explained in Section III.4. Using the SB-based machine causes the communication time between the CPU and the FPGA, although it is a small fraction of the total time. Nevertheless, the overall computation time for NBMF, including these additional factors, was shorter than that for NMF.
Throughout this study, the ratio of the number of features to the number of users was fixed at $k/n=0.5$. In general, the RMSE tends to decrease as the ratio increases. However, the computation time increases with the ratio because the matrix sizes of $H$ and $W$ increase. Therefore, a moderate value needs to be selected for this ratio. As shown in Figure 8, the rate of improvement in the RMSE was slow for ratios of 0.5 or greater across all datasets. Considering this result, we chose $k/n=0.5$ as an appropriate value for this ratio.
Our results support the computational advantages of NBMF. However, several limitations exist. Notably, the performance of NBMF is susceptible to the characteristics of datasets. NMF, which operates on continuous variables, shows comparable or superior accuracy in certain instances compared to NBMF. This higher accuracy is due to the greater resolution of continuous representations compared to binary ones. Furthermore, it is necessary to employ a low-latency system to realize the advantage of computation time. Even with a high-performance Ising machine, communication overhead between the CPU and the Ising machine can significantly impact overall performance. Therefore, utilizing a low-latency Ising machine is crucial for efficiently executing NBMF.
V Conclusions
In summary, we proposed a novel approach that employs NBMF with masking for collaborative filtering, and our findings demonstrate a substantial improvement in estimation performance as a result of the masking procedure. Moreover, our results highlighted the computational advantage of employing an SB-based machine in NBMF. NBMF with masking can be applied in collaborative filtering across various datasets. This study reveals the potential of NBMF by utilizing an Ising machine for a wide range of applications.
The efficacy of hybrid concepts using an Ising machine can extend to other algorithms, indicating significant potential for further research in this area. Similar hybrid algorithms will become increasingly active in the future as low-latency Ising machines become more advanced and popular.
Conflict of Interest Statement
K.T. is an inventor on a Japanese patent application related to this work filed by the Toshiba Corporation (no. P2019-164742, filed 10 September 2019). This study was conducted as collaborative research between Ochanomizu University and Toshiba Corporation. The authors declare that they have no other competing interests.
Author Contributions
@@ V Conclusions
Funding
This study was partially supported by JSPS KAKENHI (Grant Numbers JP23H04499 and JP25H01522) and Murata Science and Education Foundation.